### v1

# An Open-Source Annotation System for Education Transcript Data

Abstract
We introduce [Uncaptioned image] EduCoder, a domain-specialized tool designed to support utterance-level annotation of educational dialogue. While general-purpose text annotation tools for NLP and qualitative research are abound, few address the complexities of coding education dialogue transcripts—with diverse teacher-student and peer interactions. Common challenges include defining codebooks for complex pedagogical features, supporting both open-ended and categorical coding, and contextualizing utterances with external features, such as the lesson’s purpose and the pedagogical value of the instruction. EduCoder is designed to address these challenges by providing a platform for researchers and domain experts to collaboratively define complex codebooks based on observed data and incorporates both categorical and open-ended annotation types along with contextual materials. Additionally, it offers a side-by-side comparison of multiple annotators’ responses, allowing comparison and calibration of annotations with others to improve data reliability. The system is open-source [^1] https://anonymous.4open.science/r/EduCoder-5878 with demo video available [^2] https://www.youtube.com/watch?v=5YFR2seN0_k&t=4s.
EduCoder:
An Open-Source Annotation System for Education Transcript Data
Guanzhong Pan1, Mei Tan2, Hyunji Nam2, Lucía Langlois2, James Malamut2, Liliana Deonizio2, Dorottya Demszky2 1Carnegie Mellon University, 2Stanford University
1 Introduction
Teacher-student and student-student dialogue play a critical role in learning, offering insight into instruction quality, student reasoning and understanding (Mercer and Howe, 2012; Michaels et al., 2008). In recent years, the release of an increasing number of conversational datasets reflects growing interest in analyzing such dialogues (Stasaski et al., 2020; Suresh et al., 2022; Demszky and Hill, 2023; Wang et al., 2024; Macina et al., 2023). Since manual analysis is labor-intensive and difficult to scale, there has been growing efforts to develop natural language processing (NLP) based measures of educationally relevant constructs in transcript data (D’Mello and Graesser, 2012; Jensen et al., 2021; Hunkins et al., 2022; Demszky et al., 2021; Alic et al., 2022; Suresh et al., 2018; Zhao et al., 2024). However, progress in NLP model development for analyzing teachers’ and students’ talk moves hinges on rigorous and thoughtful human annotation, critical for model validation and training.
Challenges & Motivation. Rigorous human annotation of educational dialogue is nontrivial, reconciling literature-based learning theories, practitioner expertise, linguistic features, and contextual information. It follows an iterative calibration process where researchers and practitioners collaboratively specify annotation codebook. In addition to conceptual factors, there are multiple practical factors making process quite challenging.
First, no purpose-built, adaptive tool exists for annotating educational transcripts with iterative co-design and calibration support. Most existing platforms target general NLP taggings like named entity recognition (NER) or part-of-speech (POS) tagging, and lack support for annotating high-inference pedagogical features such as the discourse moves of teachers and students–and the latent cognitive processes they enact (Nakayama et al., 2018; Perry, 2021). Qualitative analysis platforms ATLAS.ti Scientific Software Development GmbH (2023); Dhakal (2022) support open coding but lack support for structured, utterance-level annotation and real-time analysis for collaborative calibration. As a result, many researchers rely on spreadsheets or ad hoc solutions, which are time-consuming to build, often clunky to use, and are poorly integrated with modern machine learning workflows.
Second, while non-technical annotators—such as educators and qualitative researchers—may possess deep domain expertise, they often face significant barriers when using computational annotation tools. These challenges extend beyond a steep learning curve to include usability issues, interface complexity, and inefficiencies in accessing dispersed contextual information. Although conventional tools like spreadsheets or manual methods may be more accessible, they often lack the integrated features necessary for high-quality analysis—such as support for cross-annotator comparison, critical for effective annotator training and reliable results. Additionally, they typically do not provide a user-friendly interface that consolidates all relevant contextual materials into a single, centralized workspace for improving annotation efficiency and supporting annotator understanding.
Third, there is growing interest in leveraging large language models (LLMs) as annotation tools to help meet the demand for high-quality labeled data. While LLMs have shown promising performance on tasks such as sentiment analysis, political stance detection, and educational content classification—sometimes achieving IRR comparable to that of humans in well-defined domains (Bojić et al., 2025; Long et al., 2024)—their reliability remains less understood when applied to more complex or socially nuanced constructs. Recent work has proposed treating LLMs as additional “raters” to help quantify and control measurement error in human annotation processes Broska et al. (2025); Calderon et al. (2025). Building on this line of research, we take it a step further: rather than treating LLMs as standalone annotators, we incorporate their outputs as one of several points of reference during human annotator calibration. This approach allows us to explore how LLMs might support, rather than replace, human judgment in complex annotation tasks.
Our System EduCoder addresses these challenges through an intuitive interface designed for utterance-level annotation in educational dialogue analysis. The system supports:
  • Customizable Codebook: Researchers can upload their own codebook of features, allowing for tailored annotation that aligns with different research questions.
  • In-text Annotation & Contextual Materials: The tool features an intuitive, side-by-side interface that enables annotators to apply codes directly to individual transcript utterances. To facilitate contextually grounded and accurate annotation, the interface presents the full transcript alongside relevant contextual materials, including task instructions and screenshots of the questions students were engaging with. Codebook references are integrated directly into the interface to provide clear definitions of features and avoid confusion along with illustrative examples and non-examples.
  • Calibration Support: EduCoder offers built-in diagnostics for IRR (Cohen, 1960), highlights disagreements to facilitate calibration, and enables comparison with other annotators to support iterative refinement of codebooks. In addition, the tool integrates annotations generated by LLMs, such as ChatGPT-4o (OpenAI, 2024) and Claude 4 Sonnet (Anthropic, 2025), to facilitate human-AI collaboration in annotation. LLM annotations can serve as an immediate reference point for human coders and support calibration. By comparing their judgments with those of other humans and LLMs, annotators can identify discrepancies, refine interpretations, and enhance their understanding of the coding task.
To demonstrate EduCoder’s flexible design and ensure its accessibility, we provide an online playground [^3] https://educoder.vercel.app/ with examples, including sample transcripts and a codebook, drawn from real-world classroom dialogue. We invite the community to use EduCoder, adapt our provided materials, and contribute to a growing ecosystem of tools for advancing educational dialogue research.
Feature EduCoder INCEpTION Prodigy Doccano Label Studio LightTag POTATO AWOCATo EASE AutoDive+
Custom Codebook import
Utterance-level annotation
Real-time IRR & visual diff
Instructions & Contextual Materials
LLM for Annotation
Usability for non-technical users
Open-source / Free for research $\times$× Free for Academia
License MIT Apache 2.0 Commercial MIT Apache 2.0 MIT Polyform Shield GPL GPL Request only
Table 1: Comparison of EduCoder with widely used annotation tools across key capabilities relevant to educational dialogue annotation. A checkmark (✓) indicates the tool fully supports the feature. A dash (–) indicates the tool provides the functionality at a basic or partial level but does not meet the specific requirements. A blank cell indicates the feature is not supported.
2 Related Works
The NLP community has benefited greatly from general-purpose annotation toolkits like INCEpTION (Klie et al., 2018), Prodigy (Montani and Honnibal, ), Label Studio (Tkachenko et al., 2020-2022), and Doccano (Nakayama et al., 2018), which standardize data labeling tasks such as NER and text classification. The newer versions of Label Studio and Prodigy also feature LLM integrations for machine coding. However, they are limited in their ability to support collaborative, iterative codebook refinement and the nuanced yet structured open-ended coding of transcripts necessary for education dialogue research. For instance, transcript annotation is often conducted at the utterance level while preserving their sequential nature—this task is ill-suited to the format of most of these tools.
Recent tools like AWOCATo (Daudert, 2020), LightTag (Perry, 2021), POTATO (Pei et al., 2022), EASE (Deng et al., 2023), AutoDive+ (Wang et al., 2025), offer more advanced features for annotation tasks to facilitate research. AWOCATo offers customizability through JSON configurations, including free-text and categorical labeling. LightTag features annotator productivity enhancements (e.g. batching, pre-annotation) and inter-annotator agreement metrics. POTATO offers extensive customization (through various task templates), active learning, and quality control mechanisms. EASE integrates multi-task active learning (utilizing annotator information) and direct integration with LLMs. AutoDive+ supports multimodal annotation, active learning, and a community-driven model repository. Despite these strengths, these tools do not adequately facilitate iterative, collaborative codebook refinement and real-time IRR metrics; nor do they offer extensive contextual referencing needed to preserve the sequential nature of transcripts and account for additional contextual information.
Qualitative Data Analysis (QDA) tools like NVivo and ATLAS.ti support interpretive coding and are well suited for qualitative education dialogue analysis. However, these tools lack support for structured utterance-level annotations essential for computational modeling and real-time IRR feedback that can facilitate annotator calibration.
Further, the integration of human judgment with AI systems has gained significant attention in NLP, particularly in reframing human-AI collaboration for explanatory tasks (Wiegreffe et al., 2021). Recent work has explored how LLMs parse and interpret text compared to human annotators (Hong et al., 2023), highlighting the importance of understanding differences in human versus AI text processing patterns (Muñoz-Ortiz et al., 2023). This research informs our approach to incorporating LLM-generated annotations as a reference while maintaining human expertise as the primary decision-making authority.
EduCoder sits at the intersection of NLP, qualitative research, and learning analytics, aiming to unify a fragmented software ecosystem. By integrating customizable codebook, real-time IRR feedback, and human-AI comparison into an accessible system, EduCoder is designed to bridge these methodological divides, supporting the multifaceted nature of educational interaction research.
3 Design Principles
EduCoder follows these core principles:
I. Minimalistic Data Structure. EduCoder requires only that the transcript include identifiable speaker and text columns, and that the codebook contain distinct code and definition columns. It also supports additional columns for usability (e.g., example columns in the codebook and segment columns indicating sections of the transcript).
II. Iterative Refinement and Data Reuse. To support the iterative nature of qualitative analysis (Weston et al., 2001; Dunivin, 2025) and the complexity of specifying nuanced social and pedagogical constructs, EduCoder allows for import and export of annotated transcripts. Researchers can load previously annotated transcripts back into the system to track how coding patterns evolve across different annotators or over time. This design enables analysis of annotation trends, distinguish between task ambiguity rooted in the codebook from systematic differences in the perspectives of individual annotators. The system incorporates established principles from case study research methodology (Eisenhardt, 1989) to support theory-building through systematic data collection and analysis.
III. Accessibility for Non-Technical Users. Usage of such computational tools should not require programming expertise to contribute domain knowledge (Park et al., 2021; Garlan et al., 2012). EduCoder provides an intuitive, checkbox-style interface with contextual reference materials and rubric definitions displayed alongside transcript views. The interface is designed to reduces cognitive load, enhance tasking understanding, and enable cross-discipline collaboration with domain experts such as educators and education researchers.
4 [Uncaptioned image] EduCoder
4.1 System Architecture
EduCoder supports two primary user roles: (1) Research Administrators who manage codebook, project settings, and export results, and (2) Annotators who perform the actual coding work. Each role has tailored interfaces that balance functionality with usability.
4.2 Annotation Workflow
Codebook Upload and Configuration. Researcher administrators can import custom codebooks that define constructs to be annotated. They can include high-level feature categories, codes, and detailed definitions with examples (and non-examples) that appear contextually during annotation to provide clear guidance to annotators in real-time and minimize labeling errors.
Transcript Import and Preprocessing. Educational transcripts can be uploaded in XLSX or CSV format with required speaker and utterance text columns, plus optional metadata such as timestamps. The system can also segment transcripts (e.g., to indicate a change in learning activity) based on a given segment column, filter by keywords/speakers and presents them in a clean interface that preserves conversational flow while enabling utterance-level annotation.
Interactive Annotation Interface. The core experience centers on a dual-pane layout with transcripts alongside feature selection panels. Annotators use checkbox interactions to mark the presence of each code, with definitions available on click.
Real-time IRR Monitoring. As annotations are completed, EduCoder computes IRR metrics both overall and per feature category, with visual indicators showing low agreement features to help identify systematic disagreements or codebook ambiguities. Such feedback accelerates the iterative codebook refinement and annotator calibration processes described by qualitative coding methodologies (Weston et al., 2001), enabling smaller batches and frequent comparison. The system implements established IRR measures, including Cohen’s kappa (Cohen, 1960) and Krippendorff’s alpha (Krippendorff, 2018), used for ensuring consistency in qualitative research (Hallgren, 2012; Hayes and Krippendorff, 2007).
4.3 Deployment and Accessibility
EduCoder can be deployed both locally or exposed publicly online. Each running instance of EduCoder serves one task with unified features. Data owners (e.g., schools or research teams) can execute private deployments for sensitive data via GitHub, while the public instance supports collaborative projects across different annotators.
The platform is released under an MIT license, ensuring broad accessibility while encouraging community contributions. We also include extensive documentation such as video tutorials, an example codebook, and step-by-step guides for common workflows. The goal is to support lowering technical barriers for educational researchers.
5 Evaluation and Case Study
To evaluate EduCoder’s usability and effectiveness, we conducted a focus group at a university summit for educators interested in the application of language technology in education. The study involved 16 K-12 mathematics teachers and two university-based math education professors. Participants completed a 45-minute training session followed by a 60-minute collaborative annotation task using common classroom dialogue transcripts. The codebook included both categorical features (e.g., "is the student playing the role of xyz?") and free-response fields capturing the annotator’s observations of students’ mathematical understanding and their relevance to the lesson’s purpose. Our evaluation methodology follows established principles for credibility in qualitative analysis (Patton, 2002) and best practices for coding in-depth interviews (Campbell et al., 2013).
The annotation process involved calibration exercises where participants annotated one to three reference transcripts to establish baseline agreement before independent coding. This small set of transcripts served as reference points for calibration during the independent annotation sessions. Annotators compared their interpretations of the codebook on the new transcripts against these reference examples to maintain consistent coding practices.
After the annotations were completed, all participants provided positive feedback about the interface and especially emphasized the ease of annotating diverse types and the value of real-time comparison with other annotators. This also prompted meaningful questions during the calibration discussions, as annotators could easily reference others’ responses and discuss whether or not they agreed with them. Teachers found the side-by-side comparison view particularly helpful for resolving disagreements and aligning on the codebook. Participants also appreciated the notes for open-ended feedback, as they could share nuanced observations that the categorical features could not fully capture.
A key challenge was calibrating different annotators’ interpretations of the open-ended codes. This is a well-known issue, often described as “complex and iterative” in qualitative research (Cole, 2024). While annotators found it easier to agree on binary labels, some of the more nuanced codes that required contextualization led to longer discussions to determine whether an observed instance represented a positive or negative example of the label. EduCoder’s visual feature highlighting the disagreement between the annotated labels were essential in identifying areas of confusion and thereby leading to targeted calibration discussions. Facilitators leveraged this feature during calibration sessions to help guide conversations. These tools encouraged participants to reflect on their own interpretations of the codes by comparing with others, and contribute to the collaborative building of the coding rubric. Several teachers suggested adding a "flag for discussion" feature to bookmark ambiguous cases for team review. The professors emphasized that surfacing subtle interpretive differences is critical for training reliable annotators and developing robust qualitative coding schemes.
Potential use-cases in teacher training and instruction reflections. The study also highlighted the potential for EduCoder to support practitioner reflection and co-design. When reviewing their own classroom conversations, teachers reported that comparing their own analysis with the LLM-generated annotations provided a valuable external perspective. This process sparked discussions about what different features of student talk the LLM was sensitive to, versus what they, as practitioners, noticed. This feedback loop suggests the tool’s utility not only for transcript annotation but also for professional development and the co-creation of research frameworks that bridge pedagogical expertise and computational analysis. The integration of LLM references address recent developments in evaluating creative and linguistic patterns in human versus AI-generated content (Ismayilzada et al., 2024), providing users with multiple perspectives on their coding decisions.
Challenges with machine annotations. However, our findings revealed that LLM annotations displayed notably low agreement with human across multiple pedagogical features. This observed challenge aligns with the prior claim Patton (2002) that coding is a complex process built on active annotator interpretation and contextual understanding. Our case study suggests that while LLMs have shown capabilities across different domains, they should not be fully trusted or used to replace expert human labelers. LLM-generated annotations may complement human annotations and serve as useful reference points for prompting calibration among annotators; however, our case study shows that they cannot fully replace the interpretative work of human annotators – especially in educational contexts, where meaningful annotations require deep pedagogical content knowledge and understanding of student’s learning trajectories.
6 Conclusion
We introduce EduCoder, an open-source annotation platform designed to enhance the study of educational dialogue. Freely available under an MIT license, it offers a user-friendly system complete with documentation, tutorials, and example materials. EduCoder incorporates best practices for qualitative analysis by supporting customizable codebooks and utterance-level coding with context. It additionally supports cross-rater comparison, incorporating LLM judgements. By providing an accessible, collaborative environment, EduCoder unifies research efforts, bridging the gap between pedagogical expertise and computational analysis to improve teaching and learning.
7 Limitations and Future Work
While EduCoder addresses several critical gaps in educational annotation tools, the current system has important limitations that constrain its applicability and suggest directions for future development.
Open-Ended Note Analysis. While the system collects free-form notes, systematic comparison across annotators remains manual and labor-intensive. Future work will implement automated analyses of open-ended notes (e.g., via topic modeling) to identify patterns in annotator responses.
Feature Category Extension. Another useful extension would be to support additional annotation types beyond categorical and free-response formats, such as Likert scales and preference labels, to better accommodate the diverse needs of annotators.

### v2

@@
 An Open-Source Annotation System for Education Transcript Data
-Guanzhong Pan1, Mei Tan2, Hyunji Nam2, Lucía Langlois2, James Malamut2, Liliana Deonizio2, Dorottya Demszky2 1Carnegie Mellon University, 2Stanford University
+Guanzhong Pan1, Mei Tan2, Hyunji Nam2, Lucía Langlois2, James Malamut2, Liliana Deonizio2, Dorottya Demszky2 1Carnegie Mellon University, 2Stanford University arthurp@andrew.cmu.edu, {mxtan, hjnam, lucilang, jmalamut, lilianas, ddemszky}@stanford.edu
 1 Introduction
@@
 First, no purpose-built, adaptive tool exists for annotating educational transcripts with iterative co-design and calibration support.
-Most existing platforms target general NLP taggings like named entity recognition (NER) or part-of-speech (POS) tagging, and lack support for annotating high-inference pedagogical features such as the discourse moves of teachers and students–and the latent cognitive processes they enact (Nakayama et al., 2018; Perry, 2021).
+Most existing platforms target general NLP taggings like named entity recognition (NER) or part-of-speech (POS) tagging, and lack support for annotating high-inference pedagogical features such as the discourse moves of teachers and students–and the latent cognitive processes they enact (Nakayama et al., 2018; Honnibal and Montani, 2017; Perry, 2021).
 Qualitative analysis platforms ATLAS.ti Scientific Software Development GmbH (2023); Dhakal (2022) support open coding but lack support for structured, utterance-level annotation and real-time analysis for collaborative calibration.